OpenStudy (anonymous):

Sloths are tree-dwelling mammals that live in South and Central America. They sleep for up to 18 hours a day. Write the portion of the day a sloth sleeps as a fraction in simplest form, and as a decimal. a. 3/4; 0.75 c. 6/8; 0.75 b. 9/12; 0.75 d. 18/24; 0.75

OpenStudy (anonymous):

We have that a sloth sleeps 18 out of the 24 hours in a day, putting this into a fraction: \[ \frac{18}{24}=\frac{18\div6}{24\div6}=\frac{3}{4}=.75 \]

OpenStudy (anonymous):

thx love

OpenStudy (anonymous):

Sure thing

OpenStudy (anonymous):

wait so it is like "a" or "d"

OpenStudy (anonymous):

Which one is in simplest terms? You tell me.

OpenStudy (anonymous):

oh a

OpenStudy (anonymous):

Yes